OpenStudy (anonymous):

Anita hikes 7 miles due east and then 3 miles due north. Rounded to the nearest mile, what is the shortest distance from the starting point to the ending point?

OpenStudy (perl):

did you try to draw it out

OpenStudy (anonymous):

a^2 + b^2 = c^2 7^2 + 3^2 = c^2 49 + 9 = c^2 58 = c^2 7.6157731058639082856614110271583 It's the closest to 8 when rounded to the nearest whole number or "nearest mile".

OpenStudy (cwrw238):